OpenStudy (anonymous):

An area of algae, A(t), grows at a rate of A'(t) = sqrt(t)*ln(t) cm^2/day. How much additional area does the algae aquire from day 1 to day 9?

OpenStudy (paxpolaris):

so you need to find:\[\Large \int\limits_{0}^{9}\sqrt t \cdot \ln(t)dt\]

OpenStudy (paxpolaris):

i think ingtegration by parts would work here

OpenStudy (paxpolaris):

\[\int\limits u dv= uv -\int\limits v du\] \[u= \ln(t) \dots du=\frac 1tdt\]\[dv= \sqrt tdt \dots v=\frac 23 t^\frac32\]

OpenStudy (anonymous):

ok thanks man.

OpenStudy (anonymous):

shouldn't it be from 1 to 9?

OpenStudy (paxpolaris):

that part i am not sure of .... i am reading the question to mean: from the beginning of the first day(t=0) to the end of the ninth day(t=9) if it means: from the END of the first day to end of ninth day, then 1 to 9 is correct

OpenStudy (anonymous):

ok thanks.